OpenStudy (dparrac123):

Solve for x and express your answer as a logarithm. 2(10^(3x) ) = 24 Simplify your answer as much as possible. Use the following notations where necessary: • For fractions, use the / symbol to separate numerator and denominator, like this: 42/53 • For logs with a base of 10, such as log107, just write the log without the base and place the value in parentheses, like this: log(7) • For logs with a base other than 10, write the base after an underscore then place the value in parentheses. For example, to write log27, write it like this: log_2(7)

OpenStudy (dparrac123):

@ganeshie8

OpenStudy (welshfella):

2 (10^3x) = 24 divide by 2 10^3x = 12 take logs log 10^3x = log12

OpenStudy (zzr0ck3r):

\(2(10^{3x} ) = 24\implies 10^{3x}=12\implies \log(10^{3x})=\log(12)\) Use \(\log_a(a^b)=b\)

OpenStudy (zzr0ck3r):

Dumb phone...Can't see when people are typing ;(